Discuss Markownikoff’s Rule and kharash effect? give  an example
 

Grade:12

3 Answers

vamshi
8 years ago
Markovnikov rule: This rule is useful when a polar reagent adds on to an unsymmetrical
alkene. It states that “The negative part of the addendum adds on to the carbon atom having less
number of hydrogen atoms”.
Mechanism of the addition of hydrogen bromide to an unsymmetrical alkene is given as follows:
HBr ionises as H+ and Br–. Electrophile, H+ attacks the double bond to form carbocation.
 
Peroxide Effect: In the presence of peroxide, addition of HBr to unsymmetrical alkenes takes place
contrary to Markovnikov’s rule. This happens only with HBr but not with HCl and HI. This
reaction is also known as Kharasch effect or Anti Markovnikov’s rule.
Peroxide
CH3-CH=CH2+HBr → CH3-CH2-CH2Br

1-Bromopropane
Propene
(major product)
It follows free radical mechanism
 
SAI SARDAR
8 years ago
Markownikoff rule:The rule states that with the addition of a protic acid HX to an unsymmetrical alkene, the acid hydrogen (H) becomes attached to the carbon with fewer alkyl substituents, and the halide (X) group becomes attached to the carbon with more alkyl substituents.
Kharash effect: kharasch effect - the process by which marchunff rule can be converted into anti marchunff rule in presence of H2O2.
Uday kiran
7 years ago
The rule states that addition of hydrogen halides to unsymmetrical alkenes takes place in such a way that the hydrogen atom(positive part of reagent) gets added to the double bonded carbon atom which has more no. of hydrogen atoms . For exampleCH3CH=CH2+HBr>CB3CHBr-CH3